OpenStudy (anonymous):

I need help on rational expression equations : 1=(1/2b)-(b+1/b)

OpenStudy (anonymous):

To solve rational equations, all you have to do is get rid of the denominators (by multiplying both sides by them), solve as you would a normal equation, and then make sure your solutions wouldn't make the original denominators equal zero. In this case, then, your first step should be to multiply both sides by \(b\), as the two fractions (\(\frac{1}{2b}\) and \(\frac{b+1}{b}\)) have the denominator \(b\).

OpenStudy (anonymous):

See if you can work it out from there.

OpenStudy (anonymous):

ok soo what we can get here is the following: 1 = (1/2b) - (b+1)/b 1 = (1/2b) - (2/2)(b+1)/b so we get 1 = (1/2b) - (2b +2)/2b 1 = (1 - 2b -2) / 2b so we get 1=-2b-1/2b 2b = -2b -1 2b+2b = -1 4b = -1 sp b = -1/4

OpenStudy (anonymous):

that one was a bit confusing. :/ i like don't the middle part D:

OpenStudy (anonymous):

well you are just trying to simplify by using common denominators because they are fractions. right? its like adding (1/2) + (1/4)